&nb

                                                                                              Normal                Real

Items Q3 2018 (billions of dollars) Gross domestic product 20,658.20           153.40

Personal consumption expenditures                              14,050.50                 111.30

Gross private domestic investment                                   3,710.70             119.40

Net exports of goods and services                                  -653.50           -108.70

Government consumption expenditures and gross investment 3,550.50    20.20

1. b) Calculate the percentage (the proportion) of each category in nominal GDP and in real GDP. Using Nominal GDP: [Personal consumption expenditures / Nominal GDP]*100% [Gross private domestic investment / Nominal GDP]*100% [Net exports of goods and services / Nominal GDP]*100% [Government consumption expenditures and gross investment/ Nominal GDP]*100% And using Real GDP: [Personal consumption expenditures / Real GDP]*100% [Gross private domestic investment / Real]*100% [Net exports of goods and services / Real GDP]*100% [Government consumption expenditures and gross investment/ Real GDP]*100% Present the information that you received in 1 (a) and 1 (b) as a table(s) in your project.

2. Write a report (2 pages double - spaced), which contains an analysis of the results you received. In this report consider, but do not be limited to the following:

1. Why was the nominal GDP greater than the real GDP? By how much?

2.GDP is composed of a number of categories. What category makes up the largest portion of GDP? What category makes up the smallest portion of GDP?

3.What is “Gross private domestic investment”? What does gross private domestic investment measure?

4.What is “Net exports of goods and services”? Why it is negative?

5.In the left part of the table 3 (nominal GDP) find the category “National defense”. How much was the National defense for the last quarter? Calculate percentage of National defense out of “Government consumption expenditures and gross investment”. Calculate percentage of National defense out of GDP.

6.Please analyze and discuss the significance of the data that you received for this Data exercise. Reflect on what you have learned from this exercise.
Answer #1

1)

At Nominal GDP

Percentage of Personal Consumption Expenditure = (Consumption Expenditure / Nominal GDP)*100 =

= (14,050.50 / 20,658.20) *100 = 68.01%

Percentage of Gross Private Domestic Investment = (3,710.70 / 20,658.20)*100 = 17.96%

Percentage of Net Exports = (653.50 / 20658.20)*100 = 3.16%

Percentage of Government consumption expenditures and gross investment = (3,550.50 / 20,658.20)*100 = 17.18%

AT REAL GDP

Percentage of Personal Consumption Expenditure = (Consumption Expenditure / Real GDP)*100 =

= (111.30 / 153.40) *100 = 72.55%

Percentage of Gross Private Domestic Investment = (119.40 / 153.40)*100 = 77.83%

Percentage of Net Exports = (108.70 / 153.40)*100 = 70.86%

Percentage of Government consumption expenditures and gross investment = (20.20/ 153.40)*100 = 13.16%

2)

A) Nominal GDP is greater than Real GDP because the former is calculated of current prices while the latter is calculated on a base year price.

Suppose, 2011 is our base year, the price of a loaf of bread in 2011 was $5 and in 2019 it is $15, so when we calculate real GDP we consider the price of bread will be taken $5 and when we consider nominal GDP price of bread will be taken as $15.

B) As per the above calculation, we can see that while considering Nominal GDP, Personal Consumption Expenditure shares the largest portion of GDP, while Export shares the smallest

In the case of Real GDP, Gross private investment shares the highest portion while the Government expenditure shares the lowest.

C)

Gross private domestic investment is the measure of physical investment used in computing GDP in the measurement of nations' economic activity. It is the amount of private business capital which is invested in domestic production either through the purchase of fixed property or inventory.

D)

Net Exports means Total amount of Goods and services exported minus Total amount of Goods and services imported.

The negative sign shows that we are importing goods and services more than exporting.